OpenStudy (anonymous):

is x^2+3x-28=0 equal to x=-7 or x-4

OpenStudy (anonymous):

x^2 + 3x- 28 = 0 x^2 +7x - 4x - 28 = 0 x(x+7) - 4(x+7) = 0 (x-4)(x+7) = 0 hence x = -4 or x= - 7